### v2

# Estimating LLM Consistency: A User Baseline vs Surrogate Metrics

Abstract
Large language models (LLMs) are prone to hallucinations and sensitive to prompt perturbations, often resulting in inconsistent or unreliable generated text. Different methods have been proposed to mitigate such hallucinations and fragility—one of them being measuring the consistency (the model’s confidence in the response, or likelihood of generating a similar response when resampled) of LLM responses. In previous work, measuring consistency often relied on the probability of a response appearing within a pool of resampled responses, or internal states or logits of responses. However, it is not yet clear how well these approaches approximate how humans perceive the consistency of LLM responses. We performed a user study ($n=2,976$) and found current methods typically do not approximate users’ perceptions of LLM consistency very well. We propose a logit-based ensemble method for estimating LLM consistency, and we show that this method matches the performance of the best-performing existing metric in estimating human ratings of LLM consistency. Our results suggest that methods of estimating LLM consistency without human evaluation are sufficiently imperfect that we suggest evaluation with human input be more broadly used.
Estimating LLM Consistency: A User Baseline vs Surrogate Metrics
Xiaoyuan Wu1. Weiran Lin1, Omer Akgul2, 1, Lujo Bauer1 1Carnegie Mellon University 2RSAC Labs Correspondence:wxyowen@cmu.edu
1 Introduction
Large language models (LLMs) have seen rapid adoption in a multitude of domains despite numerous inherent limitations such as hallucinations and fragility against adversarial attacks. Hallucinations can have disastrous consequences in high-stakes fields like healthcare and legal Merken (2025), prompting concern even as adoption continues Tiffin and Fraser (2025); Metnick et al. (2024). Further, the fragility of models enables a range of misuses that could harm users Kumar and Lakkaraju (2024); Lin et al. (2025). For instance, Lin et al. (2025) show that imperceptible changes to suggested prompts can result in outputs with biases controlled by an adversary.
Researchers have suggested that some of these issues correlate with the inconsistency of the LLMs, which is generally defined as the tendency to generate low-confidence responses or conflicting responses when the same prompt is resampled. An accurate estimation of consistency can be used: to predict if a given answer is factual, enabling the higher accuracy needed in high-stakes domains Duan et al. (2024); to detect fragile (or perhaps malicious) prompts Lin et al. (2025); in membership-inference attacks Mattern et al. (2023); to signal to users how much they should trust LLM outputs Kapoor et al. (2024); Ruggieri and Pugnana (2025).
Motivated by such utility, many have tried to define and measure the consistency of LLMs. This body of work can roughly be divided into two categories: (1) estimating consistency based on the LLMs’ internal states or logits and (2) estimating consistency based on resampling the LLMs’ responses. While the former is more compute efficient, the latter is empirically well-grounded Kuhn et al. (2023); Qiu and Miikkulainen (2024) and applicable even without white-box access to the model Lin et al. (2024).
Sampling-based estimation methods fundamentally rely on sampling from an LLM-prompt pair and comparing the outputs using a comparison function. This function differs between works—e.g., Duan et al. (2024) vs Manakul et al. (2023)—but partially boils down to estimating semantic similarity of responses. However, to our knowledge, none have based their estimation of model consistency on user-based comparisons, the ground truth for semantic similarity Bowman et al. (2015); Agirre et al. (2014). Further, the plethora of proposed uncertainty estimation methods has not yielded a clear baseline consistency metric. Without a baseline, consistency metrics are typically evaluated by their ability to predict whether a model output is factual Kuhn et al. (2023); Qiu and Miikkulainen (2024); Duan et al. (2024); Zhang et al. (2024); Kapoor et al. (2024). Joining recent work Novikova et al. (2025), we argue that consistency might be able to predict accuracy but is fundamentally an independent property of the model-prompt pair. When fed a prompt, there should be a metric that conveys to what degree the model will produce responses with equivalent meaning. We further posit that the baseline for measuring consistency is reliant on the comparison of semantics, the ground truth of which is defined by human judgement Bowman et al. (2015); Agirre et al. (2014); Nguyen et al. (2014); Raj et al. (2025).
We aim to fill this gap by using user-based comparisons to estimate the consistency of LLMs. Specifically, we conduct a user study with 2,976 participants to collect semantic similarity ratings between a sample of 10 responses to each of 100 prompts, totalling 14,880 comparisons. We then calculate response-level (one score per response) and prompt-level (one score per prompt) consistency, establishing an user-based LLM consistency baseline. Through a series of experiments, we show that existing metrics for uncertainty do not align well with human judgements collected in this study. We further show that an ensemble of logit-based scores is as similar to human judgement as the best-performing of the other methods we tested. We find that the discrepancy between existing metrics and human judgements fluctuates between models and between datasets, with previous metrics being less similar to human judgements on real-world prompts then on synthetic prompts. Based on our results we advocate for more human-based consistency evaluation in future work.
We structure the remainder of the paper in four sections: 2 details background work; 3 describes our model- and prompt-selection strategy, user study, and consistency calculation; 4 reports our experiments and comparisons; finally, we summarize and discuss implications of our work in 5.
Figure 1: Overview of our study design
2 Related Work
Here we review work in consistency estimation of LLMs with resampling and internal model states. (2.1). We also cover work that investigated logit-based metrics to estimate LLM consistency (2.2). Finally, we reiterate our study motivation (2.3).
2.1 LLM Consistency from Sampling and Internal States
Prior definitions of model consistency Lakshminarayanan et al. (2017) do not apply to the near-infinite output space, auto-regressive nature of LLMs Kuhn et al. (2023). As a result, a line of work has emerged to define uncertainty of LLM responses (sometimes defined as the confidence of the model in a response, or how likely a response is to be generated when resampled).
Sampling
When estimating uncertainty, researchers often resample multiple responses from a model-prompt pair. A chosen response is then compared to the set of responses, ultimately calculating an uncertainty score. For instance, Kuhn et al. (2023) establishes a set of semantically equivalent responses within the sampled set using entailment. The probabilities of any one of these semantically equivalent groups are fed into a predictive entropy based formulation to obtain an uncertainty score. Follow-up work has suggested similar sampling-based metrics Qiu and Miikkulainen (2024); Duan et al. (2024).
Internal state representation
Some prior work has sidestepped the definition of uncertainty to directly predict if a given answer is right or wrong. This body of work (sometimes also referred to as uncertainty estimation), uses techniques such as asking follow-ups Kadavath et al. (2022); Sam et al. (2025) and training auxiliary prediction models based on internal states Kapoor et al. (2024); Kossen et al. (2024).
2.2 Probability- and Logit-Based Uncertainty Metrics
Existing work suggests the confidence of individual token generation can suggest the consistency of LLM responses Manakul et al. (2023). Borrowing definitions from Manakul et al. (2023), we denote the probability distribution of an individual token generated as $p$, and the entropy of individual token generation $H$, $H=-\sum_{i}p_{i}log(p_{i})$. Existing metrics include the average of the minus log probability $Avg(-log(p))$, maximum of the minus log probability $Max(-log(p))$, average of entropy $Avg(H)$ and maximum of entropy $Max(H)$. In this paper, we ensemble these metrics (and variants) in 3.3.
Difference of Logits Ratio Loss
Besides probability-based (i.e., $p$-based) uncertainty metrics, we also try to find logit-based uncertainty metrics to suggest LLM response consistency. We denote the logits of individual token generation as $l$, $p=softmax(l)$. Specifically, we denote the largest, second largest, third largest, and fourth largest logit of a token generation as $l_{1}$, $l_{2}$, $l_{3}$, and $l_{4}$ correspondingly. The Difference of Logits Ratio Loss, i.e., the DLR loss, formally $\frac{l_{1}-l_{2}}{[l_{1}-(l_{3}+l_{4})]/2}$, is a loss function commonly used by adversarial attacks to suggest how confident a model is Croce and Hein (2020). In this paper, we borrow the DLR loss as a logit-based uncertainty of a token generation to estimate the consistency of LLM responses.
Since probability-based uncertainty estimation is ultimately derived from token logits, for the rest of the paper we call this group of uncertainty estimation logit-based methods.
2.3 Importance of User Evaluation in NLP Tasks
Human evaluations are important in improving and assessing the quality of natural language processing (NLP) Gritta et al. (2024); Boyd-Graber et al. (2022); Blodgett et al. (2024). Researchers have investigated how well machines are able to translate sentences into another language, compare sentences for semantic similarity, and many other NLP tasks with human evaluations Chatzikoumi (2020); Graham et al. (2017). This trend has continued in the LLM era Ouyang et al. (2022); Bai et al. (2022) While prior work has found ways to estimate how consistent responses produced by LLMs are, we argue that consistency is defined by what degree humans see responses as similar to each other. Using this definition, we establish a baseline through a user study.
3 Methods
In this study, we seek to understand how good existing metrics can approximate users’ perceptions of LLM consistency. We explore ways to create novel and more efficient ways of estimating LLM consistency without the need for expensive user studies or generating multiple responses per prompt. We provide an overview of our study design in figure 1.
In this section we first describe the prompts, models, and responses used in the study (3.1). We then describe in detail how we conducted the user study with 2,976 participants (3.2). Finally, we explain how we calculate consistency, both with existing metrics and our user-based method (3.3).
3.1 Prompts, Models, and Responses
Prompts
Studies of LLMs’ consistency fundamentally rely on a fixed set of prompts Geng et al. (2024). In this work, we selected two prompt sets: (1) To ensure comparability with prior research, we use a dataset of open-ended questions from CoQA Reddy et al. (2019), which is commonly adopted in the literature; and (2) To better represent real-world use cases, we include a sample of 50 prompts from LMSYS-Chat-1M Zheng et al. (2023).
Each entry in the CoQA dataset contains a story with a series of questions about the story. For our study, we only used the first question in each series as many following questions are dependent on the previous question or answer for context. We randomly sampled 50 story-question pairs from the CoQA dataset.
Real-world prompts from the LMSYS dataset are more diverse compared to those in CoQA. As such, we used labels in the data to filter out prompts that are non-English or Flagged111by OpenAI Moderation Zheng et al. (2023) (e.g., harassment, violence). Next, we randomly sampled 100 prompts from the filtered LMSYS dataset and manually removed 48 prompts – 14 coding questions, 9 about the LLM itself, 9 that were inappropriate (e.g., sexual), 6 that contain nonsensical sentences, 5 that we expected to have non-English responses (e.g., translations), 3 asking for answers over 1,500 words, and 2 that contain time sensitive information. We excluded coding questions since participants might not have the background to rate coding-related responses. We randomly removed 2 prompts from the remaining 52 to achieve a final sample of 50.
Models
A generalizable definition of consistency should be model-agnostic. Thus, we used three open-weights models from competeing institutions for generating responses: Llama-3.2-3B-Instruct (Llama), Gemma-2-9B-it (Gemma), and Mistral-7B-Instruct-v0.3 (Mistral). All three were the most recent publically available versions of their repective families, were widely used in prveious work, and met our hardware memory constraints (Nvidia RTX A6000). We used open-weights models as they provide us access to logits, which are necessary for our calcualtion described in section 3.3, and enables reproducibility of our study Ma et al. (2024). We left temperature settings unchanged from the default configuration because there is no consensus from pervious work on optimal temprature for uncertainty estimation Cecere et al. (2025); Du et al. (2025); Wang et al. (2023); Renze and Guven (2024); Zou et al. (2023)
Responses
The 100 prompts were randomly assigned to the three LLMs, Gemma receiving 34, Llama 33, and Mistral 33. Using the assigned model, we then generated 10 responses per prompt, totaling 1,000 responses. We took this approach as previous work has shown 10 responses adequately captures the semantic diversity of the response space for uncertainty estimation purposes Kuhn et al. (2023); Qiu and Miikkulainen (2024). Response generation took 30 minutes of GPU time.
3.2 User Study
While many metrics have been created to approximate LLM consistency, users’ perceptions of how consistent LLMs are has been largely missing from the literature. Our user-rating based approach establishes such baseline. We recruited users from Prolific and required participants 18 years or older, reside in the United States, read and type in English fluently, and have at least 95% approval rate on the platform. We use Prolific over MTurk because recent work has shown superior data quality Tang et al. (2022); Moss and Litman (2018).
Prolific users who are eligible encountered the post of our study and would be able to participate. Each participant was given an introduction about the study before seeing the instructions and examples of how to rate semantic similarity between a pair of sentences using a 6-point scale (section A.1). This commonly used 6-point scale, with explanations and examples were adopted from Agirre et al. (2014), asking crowdsourced users to rate similarity between pairs of sentences. After the instructions, each participant rated the semantic similarity of five pairs of sentences from five different prompts. Lastly, we collected participants’ demographics. The sentence pairs and their orders within the survey were randomized. Our study was approved by our institution’s ethics review board.
Model Num. Num. Num. Pairs of
Name Prompts Responses Unique Responses
Gemma 34 340 793
Llama 33 330 826
Mistral 33 330 1,319
Table 1: Breakdown of number of prompts, responses, and pairs of unique responses per model.
For each prompt, there are $\binom{num\_responses}{2}$ number of pairs of responses. For the 100 prompts, we obtained 1,000 responses, of which 756 are unique. These responses made up 4,500 pairs of responses, of which 2,938 are unique. More details about number of prompts, responses, and pairs of unique responses for each model are described in table 1.
3.3 Calculating and Comparing Consistency
Here we outline how we calculate consistency scores (we calculate similarities between pairs of responses, consistency for each response, and consistency for each prompt) and how we compare user-based scores to previous work.
User ratings of similarities
We first aggregate user ratings of similarity between a pair of responses $r_{a}$ and $r_{b}$ as the average ratings of $n\geq 5$ participants removing the highest and lowest scores Curran (2016) formally,
$s(r_{a},r_{b})=\frac{\sum\limits_{k=1}^{n-2}h_{k}}{n-2}$ (1)
Response-pair level comparison
To understand how existing metrics used for LLM consistency compare to human ratings, we compare them at the response-pair level. For each prompt (e.g., “What is Wh in batteries?”), 10 responses make up 45 response-pairs making a total of 4,500 response-pairs for the 100 prompts in our study. For these response-pairs, we calculate a Spearman correlation coefficient $\rho$ between $s_{\text{user}}$ and each of the four existing metrics—$s_{\text{\emph{Bert}}}$, $s_{\text{\emph{BLEU}}}$, $s_{\text{\emph{Rouge}}}$, and $s_{\text{\emph{USE}}}$. We use Spearman correlation coefficient $\rho$ because human ratings were on a 6-point likert scale, and we do not assume linear correlation between human ratings and existing metrics. We share our findings on how each of these compare to human ratings in section 4.2.
We calculated Bert using the BERTScore python package [^2] https://pypi.org/project/bert-score/0.3.0/, BLEU using NLTK python package [^3] https://www.nltk.org/_modules/nltk/translate/bleu_score.html, Rouge using the rouge-score python package [^4] https://pypi.org/project/rouge-score/, and USE using the universal-sentence-encoder model on Kaggle [^5] https://www.kaggle.com/models/google/universal-sentence-encoder.
Response to response-set consistency
Given that $m$ responses are generated for one prompt, using similarity scores between pairs of responses, we calculate the response to response-set consistency by averaging the similarity between one response to all other $m-1$ responses to a prompt. Formally,
$C(\text{prompt},r_{i})=\frac{\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m}s(r_{j},r_{i})}{m-1}$ (2)
In addition to the four metrics used in response-pair level comparison, we add Semantic Entropy (SEKuhn et al. (2023) at this level of comparison as their definition of consistency builds on the response to response-set level. To understand how well these five metrics perform at this level, we compute Spearman correlation coefficient and the mean squared error between each of the metrics and human ratings in section 4.3.
Additionally, motivated by reducing the overhead of resampling, we compare logits-based scores to human ratings and propose an ensembling method by using a linear combination of logits in section 4.3.
To calculate SE, we used the roberta-large-mnli model to check entailment between pairs of responses [^6] https://huggingface.co/FacebookAI/roberta-large-mnli. We provide more details on logit-based scores and our ensembling method in subsection 3.3.
Prompt level consistency
Building up from response to response-set consistency, we define prompt consistency as the average of the $m$ response to response-set consistency for all responses to the prompt, formally,
$C(\text{prompt})=\frac{\sum\limits_{i=1}^{m}C(\text{prompt},r_{i})}{m}$ (3)
We use Spearman correlation and mean squared error to determine how good each of the existing metrics and our ensemble is at approximating human ratings of prompt level consistency. We share our findings in section 4.4.
Estimating consistency with logit-based metrics.
We estimate the consistency of LLM responses using token-based uncertainty metrics, as previous work does Manakul et al. (2023). Specifically, we use four uncertainty metrics: the probability, minus log probability, entropy (all three introduced in 2.2), and the DLR loss (introduced in 2.2). For each response, we measure the maximum, sum, minimum, and average of these four token-based metrics, totaling $16$ values. The maximum and minimum suggest the ranges on individual tokens, while the average suggests the expected value across all tokens. The sum suggests a cumulative estimate over the whole response, as different responses may have different lengths. Unlike existing work that detects hallucination in one response Manakul et al. (2023), we collect several responses to estimate the consistency of LLM for a prompt, corresponding to the definition we give in equation 3.
Ensembling uncertainty metrics
In addition to individual uncertainty-based metrics described in the previous paragraph, we further perform an ensemble of these to approximate human ratings of LLM’s consistency. We used a Sequential Feature Selector scikit-learn developers (2025) to determine the most important metrics before ensemble a linear combination to approximate human ratings.
4 Results
Here we first summarize the user study data collection (4.1), the results of which will be used to establish our baseline for consistency. To compare previous consistency methods to our user-rating based method, we follow a bottom up approach. First we compare how methods of estimating similarity between sentences (or LLM responses in our case) compare to user ratings (4.2). Next, we use these similarity scores (including user-based scores) to build consistency scores for individual responses (4.3). Here we also compare logit-based uncertainty estimation methods to our user baseline. Finally we aggregate consistency scores to a set of responses to obtain a prompt-level consistency score (4.4). For each step of the hierarchy, we compare automated methods of estimating consistency to our user-rating based ground truth. We find that while some methods are closer, none are able to match the user baseline. We further find that ensembling logit-based scores approximate the best-performing sampling method.
4.1 User Study Overview and Demographics
We recruited 2,976 participants from Prolific, asking each to rate five unique response pairs from different prompts. About 52% of participants were female and 47% male. Participants’ ages ranged from 18 to 88 with various education, ethnicity, and income backgrounds. More details are provided in table 5. We collected Prolific ID for compensating participants who completes our survey. We removed Prolific ID before doing any analysis.
Model CoQA LMSYS
Prompt Reponse Prompt Reponse
Gemma 14.00 18.96 11.00 37.08
Llama 14.44 18.54 11.59 46.29
Mistral 14.69 49.18 13.85 133.83
Table 2: Average prompt and response lengths for each model and dataset.
We used 100 prompts and generated 10 responses for each prompt. Each of the 10 responses per prompts then made up 45 response pairs. Removing duplicate response-pairs (including pairs with identical responses, see section A.2), we were left with 756 unique response-pairs, each of which were rated by five or more participants.
Overall, the prompts were on average 13.33 tokens in length and the responses 52.80 tokens long. We break down the lengths for prompts drawn from the two datasets and responses generated by the three models in table 2.
4.2 Response-Pair Similarity
Our investigation starts with comparing user ratings of the similarity between response-pairs to semantic similarity metrics used in prior work. Semantic similarity metrics is a core component of sampling-based consistency metrics. We call this comparison “response-pair similarities”.
LLM Dataset Bert BLEU Rouge USE
Gemma CoQA 0.82 0.82 0.83 0.82
LMSYS 0.58 0.63 0.59 0.68
Llama CoQA 0.84 0.84 0.89 0.91
LMSYS 0.60 0.60 0.64 0.70
Mistral CoQA 0.66 0.55 0.65 0.53
LMSYS 0.57 0.57 0.52 0.51
All Both 0.71 0.74 0.73 0.75
Table 3: At the response-pair level (not consistency), USE have the highest Spearman $\rho$ correlation coefficient with human ratings overall. Additionally, existing metrics better correlate with human evaluations for prompts from the CoQA dataset than LMSYS dataset.
We evaluate the semantic similarities between each pair of responses using previously published methods and computed the correlation with human ratings. Comparing each of the four metrics evaluating the similarity of response-pairs—Bert, BLEU, Rouge, and USE Zhang et al. (2020); Papineni et al. (2001); Lin (2004); Cer et al. (2018)—to participants’ ratings, we found that no single score correlates best with how participants rate sentences’ similarities across all datasets and models. Notably, USE have the highest correlation coefficient with human ratings over all response-pairs ($\rho=0.75$) and for Llama while Bert performed the best for Mistral.
Interestingly, we found the correlation between participants’ ratings and five scores from previous work are higher for prompts from the CoQA dataset than the LMSYS dataset. Participants found response-pairs answering prompts from the CoQA dataset to be more similar to each other than response-pairs for LMSYS, by one level in the 6-point scale. To the best of our knowledge, Bert, BLEU, Rouge, and USE have not been used to evaluate LLM consistency with the more open-ended LMSYS dataset, which could have contributed to their lower correlation with human ratings.
Result 1: Among four existing metrics for evaluating the consistency of LLM responses based on similarity between pairs of sentences, USE performs the best in most cases. Additionally, we found that existing metrics better approximate human ratings on CoQA than on LMSYS.
4.3 Response to Response-Set Consistency
DLR Entropy Prob LogProb USE
Stats Mean Min. Mean Max. Mean Min. Mean Max.
Spearman $\rho$ 0.37 0.22 0.71 0.75 0.69 0.70 0.70 0.70 0.80
MSE 0.10 0.05 0.41 3.88 0.04 0.14 0.33 0.67 0.02
Table 4: At response to response-set level, logit-based scores (showing 2 best ones per type) did not perform as well as USE in approximating human ratings.
In section 4.2 we evaluted methods for measuring the similarity between pairs of responses. Consistency metrics, however, typically involve computing a score for each response that represents the extent to which that response is similar or different from other responses to the same prompt. Here we compare how a human-rating based consistency metrics performs in relation to those build with response-pair similarity metrics, as well as one additional metric that requires sampling multiple responses—SE Kuhn et al. (2023). We call this “response to response-set consistency”.
Motivated by reducing the cost of generating multiple responses to a prompt for consistency estimation, we further explore if logit-based consistency estimations methods (which have negligible cost compared to response generation) can match human-based consistency scores. For this investigation, we used 16 different logits based scores—mean, minimum, maximum, and sum for each of DLR, Entropy, Probability, and LogProbability (explained in section 3.3)—and found none of them individually approximates the consistency of LLMs as well as the USE score (see table 4).
Figure 2: At response to response-set level, our ensembling of 16 logit-based scores is as close of an approximation of human ratings as USE.
As described in section 3.3, we used an ensemble of the 16 logit-based scores to attempt to approximate human ratings, with much better success. To determine which combination of the 16 logit-based scores can create the best ensemble, we used the Sequential Feature Selector (SFS) scikit-learn developers (2025). We ran SFS 1,600 times where each number of logit-based scores (between 1 and 16) is ran 100 times. Within each of the 1,600 runs, we used 10-fold cross validation and recorded the performance (i.e., Spearman $\rho$ and MSE) of our ensemble compared to human ratings. We found using all 16 logit-based scores resulted in the highest Spearman correlation coefficient and lowest mean squared error when compared to human ratings (see figure 5). We found our ensembling method with 16 logit-based scores had a higher correlation with human ratings than Bert, BLEU, Rouge, and SE. Our ensemble, when compared to the human ratings, performed as good as USE with a 0.002 better MSE (figure 2). Same as in section 4.2, we found at response to response-set level, existign metrics correlates with human ratings better on the CoQA dataset than LMSYS figure 6.
Result 2: Using an ensemble of 16 logit-based scores can produce a consistency estimate as good as existing metrics at approximating human ratings. In addition to matching USE, our method benefits from not requiring the generation of multiple responses.
4.4 Prompt Level Consistency
As seen in the correlation difference between datasets table 3 and prior work Kuhn et al. (2023); Balabanov and Linander (2024) the prompt choice is a major contributor to consistency.
For example, responses to, “Where was the first modern Olympic Game?” is likely more consistent then responses to, “How can LLMs help users in their daily life?” As such, we next investigate consistency for a given prompt. Specifically, we aggregate response-level consistency from section 4.3 to obtain a single consistency score per prompt.
Figure 3: At per prompt level, our ensembling method and USE has the highest Spearman correlation coefficient with human evaluation of model-prompt consistency among existing metrics.
Existing metrics
Using prior defintion from section 3.3, we calculate the consistency score of each prompt using participants’ ratings and using metrics from previous work. We found that USE-based consistency scores best correlate with human-based consistency across all prompts (figure 3). We further note the consistency scores on the CoQA dataset correlates better to human ratings than those on LMSYS. This is similar to the bias we observed with pair-wise comparisons in section 4.2.
Ensembling of logit-based scores
Our ensemble of logit-based scores described in sections 3.3 and 4.3 performs as well as the USE when aggregated to the prompt level in approximating human ratings of LLM consistency. While it did not outperform USE, our method benefits from not needing to generate a set of responses and compare between responses to determine the consistency of an LLM at the prompt level.
We further investigate the number of responses needed to train an ensembled model with logit-based scores to predict the consistency of LLM. Due to the relatively small number of prompts and responses from our sample, we were not able to conclude how many responses are needed for such model to accurately predict LLMs’ consistency.
Result 3: Our ensembling method using logit-based scores performs as well as existing metrics in approximating human ratings of prompt level consistency. This approach enables estimation of LLM consistency without the need to generate multiple responses to a prompt.
5 Conclusion and Discussion
Using the human evaluation data obtained via our user study and subsequent experiments, we show that existing (automated) formulations of the consistency of LLMs are meaningfully different from our human judgment baseline. The best response to response-set consistency method achieves Spearman’s $\rho=0.80$, indicating correlation but not representation. Further, prior work has suggested that correlations considered “strong” in less precise contexts Schober et al. (2018) are not as meaningful for measuring success on NLP tasks Deutsch (2022); Bavaresco et al. (2024); Shen et al. (2023). This result holds regardless of whether we examine consistency at the per-response level (section 4.3) or at the prompt level (section 4.4).
Though sampling-based methods come closest to the human baseline, we show logit-based methods can be ensembled to approximate this performance ($\rho_{ensemble}=0.82$ vs $\rho_{\text{\emph{USE}}}=0.81$), creating an opportunity to avoid the sampling overhead.
We also find (section 4.2) that the difference in human judgement based uncertainty vs prior work is greater with real-world prompt datasets (LMSYS) compared to artificial ones (CoQA). Determining why this is the case is out of scope for our work, but we speculate that the research community’s focus thus far on artificial testing datasets might be a contributing factor.
Future directions
The discrepancy between the human baseline and existing methods raises important questions: How do imperfect consistency estimation methods affect downstream tasks? How would end-users be affected if shown consistency metrics Kapoor et al. (2024)? How are models affected when consistency measurements are part of the training cycle Liu et al. (2024)? The answers are unclear, but we believe they are worth investigating.
Finally, we urge researchers to consider the human evaluation baseline in future consistency methods revisions. We further recommend they (also) use real-world prompts for evaluation.
Limitations
Since our investigation requires logit-based scores which is often not accessible with black-box models, we used open-weights models. Such usage limits the generalizbility of our findings to black-box models. Additionally, the prompts and responses in our sample are relatively short in length (see table 2) despite us using realistic prompts from users Zheng et al. (2023), therefore, our findings may not generalize to prompts and responses of all length. Lastly, user studies are expensive to conduct. While we recuirted almost three thousand users, we were only able to evaluate 100 prompts with 10 repsonses per prompt. As we alluded to in section 5, more data from users are needed to establish a robust baseline for measuring LLM consistency.
Ethical Considerations
Our study is approved by our institution’s ethics review board. For the user study, we first provided an informed consent form to participants explaning the purpose of our survey, expected length, risks and benefits, as well as compensation. Participants who give consent to participate will proceed to the survey. As explained in section 3.1, responses shown to participants within the survey were filtered by the authors to remove any potentially harmful (e.g., harassment, sexual, violence) content so none of them were shown to any participants.
Acknowledgments
This work was supported in part by the National Institute of Standards and Technology (NIST) (ror.org/05xpvk416) and the Carnegie Mellon University (ror.org/05x2bcf33) AI Measurement Science and Engineering Center (AIMSEC); and by the PNC Center for Financial Services Innovation at Carnegie Mellon University. We thank Clement Fung for technical help with the ensemble predictor.
Appendix A Appendix
A.1 Instructions for Survey Participants
Risks
The primary risk is a breach of confidentiality since we use a third-party (Qualtrics) to design our survey and collect survey responses. Additionally, we utilize third-party vendors such as Prolific to recruit participants, and Google Drive to store and process survey responses. This risk is similar to what you encounter anytime you provide identifiable and private information online. The risks and discomfort associated with participation in this study are no greater than those ordinarily encountered in daily life or other online activities. Participants might encounter boredom or fatigue.
Figure 4: Instructions provided to participants for comparing pairs of sentences.
Instructions
A screenshot of the instructions given to participants is provided figure 4
A.2 Experiemnt asking users to rate identical responses
For each prompt, the 10 sampled responses described in section 3.1 are not all unique. We ran a 20 participant experiment on Profific with 5 pairs of unique responses for 5 different prompts. We found 17 participants gave a 5 (completely equivalent) for the identical responses. Further, three participants each gave a 4, 3, and 0 to three different pairs of identical responses. This shows participants are highly likely to rate identical pairs of responses as 5 (completely equivalent). With our approach of removing highest and lowest ratings described in section 3.3, we excluded all identical pairs of responses from the user study.
A.3 Supplemental Tables and Figures
We provide additional tables and figures here in support of our findings.
Age Num. %
18-24 290 9.74
25-34 917 30.81
35-44 708 23.79
45-54 556 18.68
55-64 324 10.89
65+ 159 5.34
Prefer not to say 22 0.74
Gender Num. %
Woman 1538 51.68
Man 1398 46.98
Other 27 0.91
Prefer not to say 13 0.44
Education Num. %
Associate’s degree 221 7.43
Bachelor’s degree 1225 41.16
Doctorate degree 137 4.60
High school graduate 298 10.01
Master’s degree 532 17.88
No high school degree 11 0.37
Professional degree 59 1.98
Some college credit, no degree 420 14.11
Trade, technical, vocational training 67 2.25
Other 1 0.03
Prefer not to say 5 0.17
Income Num. %
Under \$25K 311 10.45
\$25K to \$50K 582 19.56
\$50K to \$75K 636 21.37
\$75K to \$100K 470 15.79
\$100K or more 915 30.75
Prefer not to say 62 2.08
Table 5: Demographics data of 2,976 participants.
Table 5 provides details on demographics data of participants in our study.
Figure 5: Running 100 10-fold cross validation shows using an ensemble of all 16 logits-based scores yield the lowest Mean Squared Error and highest Spearman $\rho$ coefficient when compared to human ratings.
Figure 5 shows our procedure of selecting all 16 logit-based scores in our ensemble method.
Figure 6: At response to response-set level, existing metrics and our ensemble method correlates better with human ratings on CoQA than LMSYS.
Figure 6 shows the response to response-set level comparison between existing metrics and our ensemble method across different datasets.